OpenStudy (nuccioreggie):

Find the value of the following expression: (28 ⋅ 5−5 ⋅ 190)−2 ⋅ 5 to the power of negative 2 over 2 to the power of 3, whole to the power of 4 ⋅ 228 (5 points) Write your answer in simplified form. Show all of your steps. (5 points)

OpenStudy (anonymous):

use PEMDAS

OpenStudy (nuccioreggie):

can you explain the answer to me please

OpenStudy (nuccioreggie):

@Abhisar

OpenStudy (nuccioreggie):

@lall

OpenStudy (nuccioreggie):

@LegendarySadist

OpenStudy (nuccioreggie):

@Nnesha

OpenStudy (nuccioreggie):

@texaschic101

OpenStudy (nuccioreggie):

y u deleted ur comments

OpenStudy (anonymous):

i didnt

OpenStudy (nuccioreggie):

yea u did

OpenStudy (anonymous):

srr my computer keeps freezing

OpenStudy (anonymous):

\[(28 \times 5-5 \times 190)^{-2}\]

OpenStudy (anonymous):

\[(\frac{ 5^{-2} }{ 2^{3} })^{4}\]

OpenStudy (nuccioreggie):

so can you tell me what the anser for part A and part B will be and plexplain please

OpenStudy (anonymous):

what is 228?

OpenStudy (nuccioreggie):

can you jus tell me the answer and explain them please and i will medal you

OpenStudy (anonymous):

\[(28 \times 5−5 \times 190)^{-2} \times (\frac{ 5^{-3}}{ 2^{3} })^{4} \times 228 \]

OpenStudy (anonymous):

i am not here for medal

OpenStudy (anonymous):

now first u solve what ever is in bracket

OpenStudy (anonymous):

multiply first then subtract can u plz do that

OpenStudy (anonymous):

i am refering to the first bracket

OpenStudy (anonymous):

\[(-810)^{-2} \times (\frac{ 1 }{ 5^{2} \times 2^{3} })^{4} \times 228\]

OpenStudy (anonymous):

\[(-810)^{-2} \times (\frac{ 1^{4} }{ 200^{4}}) \times 228\]

OpenStudy (nuccioreggie):

so what the aswer for part A?

OpenStudy (anonymous):

\[-\frac{ 1 }{ 810^{2} } \times \frac{ 1 }{ 200^{4} } \times 228\]

OpenStudy (anonymous):

where is the part A

OpenStudy (anonymous):

there is no part A in the question just say solve

OpenStudy (nuccioreggie):

O YEA SORRY

OpenStudy (nuccioreggie):

LOL whats the answer for this question can you please explain

OpenStudy (anonymous):

\[-\frac{ 1 }{ 656100} \times \frac{ 1 }{ 1600000000 } \times 228\]

OpenStudy (anonymous):

cant u seem to follow the steps i am doing

OpenStudy (anonymous):

u need to write all these steps to get 5 points

OpenStudy (nuccioreggie):

i need to hurry up and get the answer because this has a time limit i have 5 more minutes thats y i just asked for the answer so please can you give me the answer and explain it pplease

OpenStudy (anonymous):

p Parentheses first E Exponents (ie Powers and Square Roots, etc.) MD Multiplication and Division (left-to-right) AS Addition and Subtraction (left-to-right)

OpenStudy (anonymous):

use a calculator and calculate , i have given u everything

OpenStudy (anonymous):

@imqwerty does my solution makes sense?

imqwerty (imqwerty):

well what u did is all correct but usually we get nice solutions nd not like that we r getting here maybe the question is not interpreted well... but if what u wrote is the question then the solution is correct :)

OpenStudy (anonymous):

@imqwerty yea i was wondering the same thing

imqwerty (imqwerty):

:)